# Relaxation of quasi-convex functionals with variable exponent growth

### v1

@@ 1. Introduction and main result
Existence theory of several models have been considered, related to optimal design and double phase problems, thin structures, dielectric materials, homogenization and fluids: we mention [5, 6, 7, 8, 11, 17, 18, 27, 35], among a wider literature, where lower semicontinuity, relaxation and variational convergence for integral functionals with densities satisfying a $p(\cdot)$}-growth condition have been considered, under a variety of continuity assumptions on the exponent pp. In these papers integral representations are obtained, more or less explicitly, according to the regularity hypotheses on pp.
@@
The main contribution of this paper is to introduce the first model which features both quasi-convex anisotropy of [3, 20, 21] and the combination of linear and super-linear growth of [16, 24, 25]. We provide sufficient conditions for lower semicontinuity and relaxation of functionals of the form (1.1) when the exponent $p$ may take the value 11. Once again, the $\log$-Hölder continuity ensures the lack of the Lavrentiev phenomenon and that the only concentration effect is the expected one from the BV-space described by the recession function. However, the subtlety of the problem is illustrated by the fact that a slightly stronger vanishing $\log$-Hölder continuity is required in the linear growth set $\{p=1\}$.
@@
Our proof is based on approximation by linear growth functionals, for which we can use an improved version of the linear result due to Breit, Diening and Gmeinder [9], see Lemma 2.4. Based on this lemma we obtain in Proposition 4.2 a lower semi-continuity result which allows us to prove the inequality $\geqslant$” in the main theorem. The proof of the opposite inequality uses the so-called blow-up method from [4, Section 5.5], which, as a first step, requires us to prove that $\mathcal{F}(u,\cdot)$ in (1.2) is a Radon measure absolutely continuous with respect to the generalized total variation of DuDu, in the sense of [16, 24, 25]. In Proposition 4.4 we refine techniques from [25] to show how the strong $\log$-Hölder continuity allows us to separate the effect of the exponent into a multiplicative constant which tends to 11 in the blow-up process. The remaining linear growth, quasi-convex part is then handled as in [4, Section 5.5]. Before the main result, we introduce background material in Section 2 and the appropriate vector-valued BV-type space with variable growth in Section 3.
@@ 2. Preliminaries
The variable exponent Sobolev space $W^{1,p(\cdot)}(\Omega)$ consists of functions $u\in L^{p(\cdot)}(\Omega)$ whose distributional gradient DuDu belongs to $L^{p(\cdot)}(\Omega)$. The variable exponent Sobolev space $W^{1,p(\cdot)}(\Omega)$ is a Banach space with the norm
@@
The norm of the latter term is the $L^{p(\cdot)}$-norm of the scalar-valued function $|Du|$, the Euclidean norm of DuDu, i.e. $\|Du\|_{L^{p(\cdot)}(\Omega)}:=\big\||Du|\big\|_{L^{p(\cdot)}(\Omega)}$.
@@ 3. BV-type spaces
Such functions have distributional first derivatives DuDu which are Radon measures [4, Section 3.1] and $V(u,\Omega)$ equals the total variation $|Du|(\Omega)$ of the measure DuDu [4, Proposition 3.6].
@@
Lemma 3.5 (BVBV-type approximation by smooth functions).
@@
Since $u\in W^{1,1}_{\rm loc}(\Omega;{\mathbb{R}}^{m})$, it follows from the definition of $V^{m}_{p(\cdot)}$ and integration by parts that
(3.7) $V^{m}_{p(\cdot)}(u)=\sup\bigg\{\int_{\Omega}\sum_{i=1}^{m}\nabla u_{i}\cdot w_{i}\,dx\colon w\in[C^{1}_{0}(\Omega;\mathbb{R}^{n})]^{m},\|w\|_{\varphi^{*}}\leqslant 1\bigg\},$
@@
Taking the supremum over $w\in[C^{1}_{0}(\Omega;\mathbb{R}^{n})]^{m}$ with $\|w\|_{L^{\varphi^{*}}(\Omega)}\leqslant 1$, we conclude that $V^{m}_{p(\cdot)}(u)\leqslant\|\nabla u\|_{(L^{\varphi^{*}}(\Omega))^{\prime}}$.
To prove claim (1), we next show the opposite inequality $\|\nabla u\|_{(L^{\varphi^{*}}(\Omega;{\mathbb{R}}^{m}))^{\prime}}\leqslant V^{m}_{p(\cdot)}(u)$ when smooth functions are dense. Let $w\in L^{\varphi^{*}}(\Omega;\mathbb{R}^{m\times n})$ with $\|w\|_{\varphi^{*}}=1$ and let $(w_{j})$ be a sequence from $C^{1}_{0}(\Omega;\mathbb{R}^{m\times n})$ with $w_{j}\to w$ in $L^{\varphi^{*}}(\Omega;\mathbb{R}^{m\times n})$ and almost everywhere. Since also $w_{j}/\|w_{j}\|_{\varphi^{*}}\to w$ in $L^{\varphi^{*}}(\Omega;\mathbb{R}^{m\times n})$, we may assume that $\|w_{j}\|_{\varphi^{*}}=1$. By Fatou’s Lemma,
@@
$V^{m}_{p(\cdot)}(u)\geqslant\sup\bigg\{\int_{\Omega}\sum_{i=1}^{m}\nabla u_{i}\cdot w_{i}\,dx\,\Big|\,w\in L^{\varphi^{*}}(\Omega;\mathbb{R}^{m\times n}),\|w\|_{\varphi^{*}}\leqslant 1\bigg\}.$
@@
$V^{m}_{p(\cdot)}(u)\geqslant\sup\bigg\{\int_{\Omega}|\nabla u|\,h\,dx\,\Big|\,h\in L^{\varphi^{*}}(\Omega),\|h\|_{\varphi^{*}}\leqslant 1\bigg\}=\|\nabla u\|_{(L^{\varphi^{*}}(\Omega))^{\prime}}.$
Hence $V^{m}_{p(\cdot)}(u)=\|\nabla u\|_{(L^{\varphi^{*}}(\Omega))^{\prime}}$ and the proof of (1) is complete.
@@
$V^{m}_{p(\cdot)}(u)\geqslant\sup\bigg\{\int_{\Omega}|\nabla u|\,h\,dx\,\Big|\,h\in C^{1}_{0}(\Omega),\|h\|_{\varphi^{*}}\leqslant 1\bigg\}.$
@@
In [24, 25], a BV-type space with variable exponent was defined based on the set $Y$ where the exponent equals 11 using the modular which is denoted $\varrho_{\text{old}}$ below. The next result connects this definition to the one given above based on [16].
@@
where $p^{\prime}$ is the Hölder conjugate of pp. Since $p$ is lower semicontinuous, $Y=\{p\leqslant 1\}$ is closed. Since further $u\in W^{1,p(\cdot)}(\Omega\setminus Y)$, $D^{s}u=0$ in the open set $\Omega\setminus Y$. Then $|w|\leqslant 1$ and Young’s inequality imply that
@@ 4. Proof of main result
If $u\in{\rm BV}^{p(\cdot)}(\Omega;{\mathbb{R}}^{m})$, then the previously shown inequality ensures that $\mathcal{F}(u,A)$ is finite; since $f\geqslant 0$, it is increasing and non-negative. The superadditivity condition (ii) is a consequence of the definition of $\mathcal{F}$ since $u_{h}\in W^{1,{p(\cdot)}}(A\cup A^{\prime};{\mathbb{R}}^{m})$ if and only if $u_{h}|_{A}\in W^{1,{p(\cdot)}}(A;{\mathbb{R}}^{m})$ and $u_{h}|_{A^{\prime}}\in W^{1,{p(\cdot)}}(A^{\prime};{\mathbb{R}}^{m})$ for disjoint and open AA and $A^{\prime}$.
@@
Since $p^{+}<\infty$, the modular $\varrho_{\text{\rm old}}(u)$ over AA is finite, and hence from the definition of $\varrho_{\text{\rm old}}$ we obtain that $\|u\|_{\varrho_{\text{\rm old}},F}\to 0$ as $F\to\emptyset$. Since AA is open, $A\setminus\bar{A}_{2}\to\emptyset$ as $A_{2}\nearrow A$. Then (iii) follows by $\mathcal{F}(u,A)\leqslant\mathcal{F}(u,A_{1})+\mathcal{F}(u,A\setminus\overline{A}_{2})$ as $A_{2}\nearrow A$.
To show (i), we replace AA by $A\cup A^{\prime}$ in (4.6), choose $A_{1}\subset\subset A$ and use monotonicity:

### v2

@@ 1. Introduction and main result
Existence theory of several models have been considered, related to optimal design and double phase problems, thin structures, dielectric materials, homogenization and fluids: we mention [5, 6, 7, 8, 11, 17, 18, 27, 35], among a wider literature, where lower semicontinuity, relaxation and variational convergence for integral functionals with densities satisfying a $p(\cdot)$}-growth condition have been considered, under a variety of continuity assumptions on the exponent $p$. In these papers integral representations are obtained, more or less explicitly, according to the regularity hypotheses on $p$.
@@
The main contribution of this paper is to introduce the first model which features both quasi-convex anisotropy of [3, 20, 21] and the combination of linear and super-linear growth of [16, 24, 25]. We provide sufficient conditions for lower semicontinuity and relaxation of functionals of the form (1.1) when the exponent $p$ may take the value $1$. Once again, the $\log$-Hölder continuity ensures the lack of the Lavrentiev phenomenon and that the only concentration effect is the expected one from the BV-space described by the recession function. However, the subtlety of the problem is illustrated by the fact that a slightly stronger vanishing $\log$-Hölder continuity is required in the linear growth set $\{p=1\}$.
@@
Our proof is based on approximation by linear growth functionals, for which we can use an improved version of the linear result due to Breit, Diening and Gmeinder [9], see Lemma 2.4. Based on this lemma we obtain in Proposition 4.2 a lower semi-continuity result which allows us to prove the inequality $\geqslant$” in the main theorem. The proof of the opposite inequality uses the so-called blow-up method from [4, Section 5.5], which, as a first step, requires us to prove that $\mathcal{F}(u,\cdot)$ in (1.2) is a Radon measure absolutely continuous with respect to the generalized total variation of $Du$, in the sense of [16, 24, 25]. In Proposition 4.4 we refine techniques from [25] to show how the strong $\log$-Hölder continuity allows us to separate the effect of the exponent into a multiplicative constant which tends to $1$ in the blow-up process. The remaining linear growth, quasi-convex part is then handled as in [4, Section 5.5]. Before the main result, we introduce background material in Section 2 and the appropriate vector-valued BV-type space with variable growth in Section 3.
@@ 2. Preliminaries
The variable exponent Sobolev space $W^{1,p(\cdot)}(\Omega)$ consists of functions $u\in L^{p(\cdot)}(\Omega)$ whose distributional gradient $Du$ belongs to $L^{p(\cdot)}(\Omega)$. The variable exponent Sobolev space $W^{1,p(\cdot)}(\Omega)$ is a Banach space with the norm
@@
The norm of the latter term is the $L^{p(\cdot)}$-norm of the scalar-valued function $|Du|$, the Euclidean norm of $Du$, i.e. $\|Du\|_{L^{p(\cdot)}(\Omega)}:=\big\||Du|\big\|_{L^{p(\cdot)}(\Omega)}$.
@@ 3. BV-type spaces
Such functions have distributional first derivatives $Du$ which are Radon measures [4, Section 3.1] and $V(u,\Omega)$ equals the total variation $|Du|(\Omega)$ of the measure $Du$ [4, Proposition 3.6].
@@
Lemma 3.5 ($BV$-type approximation by smooth functions).
@@
Since $u\in W^{1,1}_{\rm loc}(\Omega;{\mathbb{R}}^{m})$, it follows from the definition of $V^{m}_{\varphi}$ and integration by parts that
(3.7) $V^{m}_{\varphi}(u)=\sup\bigg\{\int_{\Omega}\sum_{i=1}^{m}\nabla u_{i}\cdot w_{i}\,dx\colon w\in[C^{1}_{0}(\Omega;\mathbb{R}^{n})]^{m},\|w\|_{\varphi^{*}}\leqslant 1\bigg\},$
@@
Taking the supremum over $w\in[C^{1}_{0}(\Omega;\mathbb{R}^{n})]^{m}$ with $\|w\|_{L^{\varphi^{*}}(\Omega)}\leqslant 1$, we conclude that $V^{m}_{\varphi}(u)\leqslant\|\nabla u\|_{(L^{\varphi^{*}}(\Omega))^{\prime}}$.
To prove claim (1), we next show the opposite inequality $\|\nabla u\|_{(L^{\varphi^{*}}(\Omega;{\mathbb{R}}^{m}))^{\prime}}\leqslant V^{m}_{\varphi}(u)$ when smooth functions are dense. Let $w\in L^{\varphi^{*}}(\Omega;\mathbb{R}^{m\times n})$ with $\|w\|_{\varphi^{*}}=1$ and let $(w_{j})$ be a sequence from $C^{1}_{0}(\Omega;\mathbb{R}^{m\times n})$ with $w_{j}\to w$ in $L^{\varphi^{*}}(\Omega;\mathbb{R}^{m\times n})$ and almost everywhere. Since also $w_{j}/\|w_{j}\|_{\varphi^{*}}\to w$ in $L^{\varphi^{*}}(\Omega;\mathbb{R}^{m\times n})$, we may assume that $\|w_{j}\|_{\varphi^{*}}=1$. By Fatou’s Lemma,
@@
$V^{m}_{\varphi}(u)\geqslant\sup\bigg\{\int_{\Omega}\sum_{i=1}^{m}\nabla u_{i}\cdot w_{i}\,dx\,\Big|\,w\in L^{\varphi^{*}}(\Omega;\mathbb{R}^{m\times n}),\|w\|_{\varphi^{*}}\leqslant 1\bigg\}.$
@@
$V^{m}_{\varphi}(u)\geqslant\sup\bigg\{\int_{\Omega}|\nabla u|\,h\,dx\,\Big|\,h\in L^{\varphi^{*}}(\Omega),\|h\|_{\varphi^{*}}\leqslant 1\bigg\}=\|\nabla u\|_{(L^{\varphi^{*}}(\Omega))^{\prime}}.$
Hence $V^{m}_{\varphi}(u)=\|\nabla u\|_{(L^{\varphi^{*}}(\Omega))^{\prime}}$ and the proof of (1) is complete.
@@
$V^{m}_{\varphi}(u)\geqslant\sup\bigg\{\int_{\Omega}|\nabla u|\,h\,dx\,\Big|\,h\in C^{1}_{0}(\Omega),\|h\|_{\varphi^{*}}\leqslant 1\bigg\}.$
@@
In [24, 25], a BV-type space with variable exponent was defined based on the set $Y$ where the exponent equals $1$ using the modular which is denoted $\varrho_{\text{old}}$ below. The next result connects this definition to the one given above based on [16].
@@
where $p^{\prime}$ is the Hölder conjugate of $p$. Since $p$ is lower semicontinuous, $Y=\{p\leqslant 1\}$ is closed. Since further $u\in W^{1,p(\cdot)}(\Omega\setminus Y)$, $D^{s}u=0$ in the open set $\Omega\setminus Y$. Then $|w|\leqslant 1$ and Young’s inequality imply that
@@ 4. Proof of main result
If $u\in{\rm BV}^{p(\cdot)}(\Omega;{\mathbb{R}}^{m})$, then the previously shown inequality ensures that $\mathcal{F}(u,A)$ is finite; since $f\geqslant 0$, it is increasing and non-negative. The superadditivity condition (ii) is a consequence of the definition of $\mathcal{F}$ since $u_{h}\in W^{1,{p(\cdot)}}(A\cup A^{\prime};{\mathbb{R}}^{m})$ if and only if $u_{h}|_{A}\in W^{1,{p(\cdot)}}(A;{\mathbb{R}}^{m})$ and $u_{h}|_{A^{\prime}}\in W^{1,{p(\cdot)}}(A^{\prime};{\mathbb{R}}^{m})$ for disjoint and open $A$ and $A^{\prime}$.
@@
Since $p^{+}<\infty$, the modular $\varrho_{\text{\rm old}}(u)$ over $A$ is finite, and hence from the definition of $\varrho_{\text{\rm old}}$ we obtain that $\|u\|_{\varrho_{\text{\rm old}},F}\to 0$ as $F\to\emptyset$. Since $A$ is open, $A\setminus\bar{A}_{2}\to\emptyset$ as $A_{2}\nearrow A$. Then (iii) follows by $\mathcal{F}(u,A)\leqslant\mathcal{F}(u,A_{1})+\mathcal{F}(u,A\setminus\overline{A}_{2})$ as $A_{2}\nearrow A$.
To show (i), we replace $A$ by $A\cup A^{\prime}$ in (4.6), choose $A_{1}\subset\subset A$ and use monotonicity: